OpenStudy (anonymous):

Shelia solved this problem. A local theater sold 624 tickets for a recent performance. If the tickets were $3.00 each, how much money did it collect? Shelia said $208. Is Shelia's answer reasonable?

OpenStudy (anonymous):

A. Yes, because 600 ÷ 3 = 200, and 624 tickets were sold, so the answer will be a little more than this estimate. B. Yes, because 630 ÷ 3 = 210, and 624 tickets were sold, so the answer will be a little less than this estimate. C. No, because 600 • 3 = 1800, and 624 tickets were sold, so the answer will be a little less than this estimate. D. No, because 600 • 3 = 1800, and 624 tickets were sold, so the answer will be a little more than this estimate.

OpenStudy (dayakar):

624* 3 =1812

OpenStudy (dayakar):

D

OpenStudy (anonymous):

thankyou that helps me understand